Select the set of equations that represents the following situation: The sum of two numbers is 73, and their difference is thirty-two less than the smaller number. What are the two numbers?
I am really unclear as to this problem, I think I start it off like this,

x+y=73
But I am at a loss for the rest. Any assistance would be helpful.
Thanks
Phil

Answer by mananth(16946)

x+y=73..............1
x-y = y-32
x-2y=-32 .........2
1 x + 1 y = 73 .............1
1 x + -2 y = -32 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
2x+2y=146
x-2y=-32
Add the two equations
3x=114
/3
x=38
plug value of x in (1)
x+y=73
38+y=73
y=73-38
y=35